What is the molar mass of a .810 g sample of gas that occupies a volume of .600 L at a pressure of .777 atm and a temperature of 19.0 C? Answer in units of g/mol.
1
Expert's answer
2020-04-28T12:25:53-0400

M=m/n

n=PV/RT=(0.77*0.6)/(0.082*292)=0.462/23.944=0.019mol


M=0.810/0.019=43g/mol


Answer: m=43g/mol